Hi. Use the given zero to find the remaining zeros of each function.
h(x)=x^4-9x^3+21x^2+21x-130 Zero: 3-2i
So I know that we have 2 zeros now: (x-(3-2i)(x-(3+2i). But how can I proceed? I don't know how to divide the product of those zeros in a synthetic division.
Many thanks

Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

 
Hi,
h(x)=x^4-9x^3+21x^2+21x-130
graphing calculator/software or Trial & Error would give the remaining real roots:
x = -2 and x = 5
FREE graph software available at http://www.padowan.dk.com

Answer by richwmiller(17219)

(x^2-3x-10)(x^2-6x+13)
x^2-3x-10= (x+2) (x-5)
(x+2)(x-5)(x^2-6x+13)
(x-(3-2i)(x-(3+2i)=(x^2-6x+13)
the zeros must be factors of -130
the prime factors are -2󬊁3
There are only a few to test
-2, 2, -5 and 5 , -13 and 13